ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 90D12 Victor V. Zakharov, Saint-Petersburg University, Russia Subcore and reduced cooperative games The subcore is a multiple selector of the core in balanced TU-cooperetive games. Analitical representation of the subcore is based on an optimal solution of a linear programming problem which is used as a status quo point. A reduced cooperative game is a game with removed player (or a coalition of players). We will present a new form of reduced game and discuss an application of linear programming approach to prove consistency of the subcore in this game and also in some other reduced games. Part of these results is joint work with O-Hun Kwon. Partly supported by RBRF, project 97-01-01177. Reference: V. Zakharov, and O-Hun Kwon, Linear programming approach in cooperative games, J. of Korean Math. Soc. 34, 2 (1997), 423–435.